# A Model Zoo on Phase Transitions in Neural Networks

### v1

@@ 5 Model Zoo Applications in Weight Space Learning
Model Training.
@@
Model Analysis.
@@
Model Generation.
Finally, there is a growing push to generate NN weights, either to initialize new models without full retraining or to combine multiple networks for improved performance (Schürholt et al., 2022a; Peebles et al., 2022; Schürholt et al., 2024; Wang et al., 2024; Putterman et al., 2024; Soro et al., 2025). However, as we argue above, existing datasets to that end either contain only well-trained models or do not control for diversity in a quantifiable way. Our dataset provides precisely the sort of phase-diverse, architecture-diverse examples that generative models can learn from. By conditioning on quantifiable quality metrics that describe the phase, or on loss landscape information like curvature or connectivity, a generator trained on our zoos could be trained to produce custom weights with predictable properties.
@@ Conclusion
Our work introduces a dataset of neural networks that systematically spans multiple phases of training across different architectures and domains. By providing controlled diversity in the weight space, we aim to accelerate research on weight space learning, a growing field that treats trained weights as a data modality. This dataset can help researchers analyze when and how different training phases arise, develop phase-aware training and downstream methods, and gain deeper insights into the fundamental behavior of neural networks.
@@
KS, LM and DB are supported by the Swiss National Science Foundation research project grant 10001118. YZ and YY are supported by DOE under Award Number DE-SC0025584 and funding from Dartmouth College.
@@ Appendix A Dataset documentation
A sample of the phase transition model zoo dataset can be downloaded anonymously from https://drive.proton.me/urls/V2E66KY0JM#Pq5M06URN4EN. Further instructions to explore, visualize, or use the zoos can be found in the supplementary material and the corresponding README.md.
@@ A.2 Model Zoo Generation
We generated the dataset for common computer vision, language, and SciML tasks and architectures to show generalization and maximize applicability to the community. We fixed the load-temperature grids by exploring the boundary cases first and establishing the presence of phase transitions, then filling in more resolution. For the vision and language zoos, we chose batch size as a temperature-like parameter, as it’s a common hyperparameter to vary. For the load-like parameter, we vary the width of the models by changing the number of channels or the size of the internal token, respectively. For the SciML zoo, we adapt the scheme to the specifics of the domain. To vary the temperature, we chose to change the learning rate to keep the batch size and collocation points constant. To vary the load, we change the complexity of the domain by varying $\beta$ parameter of the physical system. We base the GPT-model implementation and training setup on NanoGPT Karpathy (2025) and use Chinchilla scaling laws to find the ideal number of training steps Hoffmann et al. (2022). The full list of model zoo hyperparameters is given in Tables 2,3 and 4. For vision zoos, we use Random Cropping, horizontal flipping, and random rotations. Training ViTs on CIFAR100 required stronger data augmentation to achieve competitive performance. Therefore, we have applied a combination of random cropping, random erasing, color jitter, and RandAugment (Cubuk et al., 2019). After the initial tuning of the grids, the training of the model zoos was done on 16 DGX H100 GPUS in 30 days. The computation of loss landscape metrics was performed on the same hardware in 14 days.
@@
        Learning Rate         $1.0$
@@ A.3 Model zoo evaluation
In this section, we test the general validity of the trained models as representatives of real-world models in a structured dataset. An overview of the models at the end of training is given in Tables 5 and 7. The results confirm that models are trained to competitive performance for their respective sizes. More nuanced information on the distribution of model performance on the temperature-load grid is shown in Figures 6 through 16. Similar to previous work, the zoos show distinct low train-loss regions, with smaller embedded regions within that generalize well. Test performance generally improves with decreasing load (increasing width), with a distinct peak phase where temperature and load are low enough, but not too low. The generalization gap correspondingly shows a superposition of both patterns. Further applications or loss landscape metrics likewise show clear phase transitions. Remarkably, the phase layout and loss landscape metrics generalize across different domains, tasks, architectures, and datasets.
@@ A.5 Hosting, Licensing, and Maintenance Plan
The dataset will be made publicly available and licensed under the Creative Commons Attribution 4.0 International license (CC-BY 4.0). The dataset will be published after the completion of the review process, for the camera-ready version. At the stage of submission, we provide representative samples of the dataset to the reviewers and will incorporate reviewer feedback for the published version of the dataset. The dataset sample is available to the reviewers at https://drive.proton.me/urls/V2E66KY0JM#Pq5M06URN4EN.
We are currently setting up a hosting plan with the University of St. Gallen to ensure long-term availability. We will collect information on the dataset, links to the data, code, and projects with the dataset on a hosted website. Upon publication, code to use the datasets will be made available, including multi-purpose dataset metadata like croissant. We will maintain the datasets if necessary. We further plan on extending the dataset towards more architectures, tasks, and domains, and invite the community to engage.
Code to recreate, correct, adapt, or extend the datasets will be provided in a GitHub repository, such that maintenance can be taken over by the community at need. The GitHub repository allows the community to discuss, interact, add, change, or fork the code.

### v2

@@ 5 Model Zoo Applications in Weight Space Learning
5.1 Learning Dynamics Analysis
In this paper, we focus on the phase transition phenomenon as studied in the statistical physics literature (Engel, 2001). This line of work links physics models, such as the Curie-Weiss model of magnetization, to problems in statistical inference. It often uses strong analytical tools like the replica method to compute free energy and characterize the phase transitions of learning systems. For a comprehensive overview, see Zdeborová and Krzakala (2016). At the same time, there is a growing body of research on phase transitions in learning dynamics. These studies show that the behavior of learning dynamics/trajectories can shift sharply as time evolves or as hyperparameters change. Some of this work also connects to statistical physics, while others develop independently. For example, Lewkowycz et al. (2020) identify three learning phases under different initial learning rates: lazy, catapult, and divergent. They show that a large but non-divergent learning rate yields the best test performance. As another example, Baity-Jesi et al. (2018) describe three stages of training: exploration of the loss landscape, aging dynamics with an increasing number of flat directions, and a final phase resembling diffusion. Extending our analysis to capture more of these phenomena in various learning dynamics is an important direction for future work.
5.2 Model Training
@@
We demonstrate the potential of our dataset for phase-aware training algorithms with a simple hyper-parameter optimization experiment to maximize performance gains in a single step. We consider any state on the grid, and attempt to change temperature or load like parameters to improve performance. That is, for each configuration, we perform one-step hyperparameter optimization (to any other load-temperature cell) and measure the resulting performance improvement. As stand-in for validation-based methods, we use a random search that keeps the post-hoc better cell in validation metrics. We compare against a simple decision rule based on the current loss landscape metrics.
Random Search Baseline: For each model configuration, we randomly select one of three tuning actions: increase model width, increase batch size, or decrease batch size. The magnitude of adjustment is also randomly sampled from a predefined range 1-5.
Loss Landscape-Guided Search: For each model configuration, we first classify its phase using four loss landscape metrics and the phase boundaries by identifying thresholds similarly to Yang et al. (2021). We then apply deterministic tuning directions:
  • Phase I and III: Increase model width
  • Phase II: Increase batch size
  • Phase IV-A: Decrease batch size
  • Phase IV-B: No parameter adjustment
While tuning directions are deterministic based on phase classification, tuning magnitudes are randomly sampled to maintain a fair comparison with the baseline. Please note that this one-step procedure can easily be extended to multi step optimization. We focus on single step solely for evaluation purposes. To ensure statistical reliability, we repeat each experiment 50 times and report both mean performance gains and standard deviations. We report results in Table 1.
Table 1: Evaluating the performance of phase-aware hyperparameters optimization approaches compared to a random search baseline, on ResNet-18 models. We report average over 50 runs as well as standard deviation. For all three model zoos evaluated, phase-aware hyperparameters optimization outperforms the baseline.
Algorithm CIFAR-10 CIFAR-100 TinyImageNet
Random Search $012\pm 0.006 0.0$$26\pm 0.009 0.0$$037\pm 0.012$
Phase-Aware Optimization $\mathbf{0.038\pm 0.004}$ $\mathbf{0.072\pm 0.002}$ $\mathbf{0.116\pm 0.007}$
The results clearly show that using loss landscape information to optimize model hyperparameters significantly improves performance and efficiency of a single optimization step. We would further like to refer the reader to Zhou et al. (2024), who have systematically evaluated the directional benefit of loss landscape metrics for hyper-parameter optimization over conventional validation metrics.
5.3 Model Analysis
@@ 5.4 Model Editing
5.5 Model Generation
Finally, there is a growing push to generate NN weights, either to initialize new models without full retraining or to combine multiple networks for improved performance (Schürholt et al., 2022a; Peebles et al., 2022; Schürholt et al., 2024; Wang et al., 2024; Putterman et al., 2024; Soro et al., 2025; Meynent et al., 2025; Falk et al., 2025b). However, as we argue above, existing datasets to that end either contain only well-trained models or do not control for diversity in a quantifiable way. Our dataset provides precisely the sort of phase-diverse, architecture-diverse examples that generative models can learn from. By conditioning on quantifiable quality metrics that describe the phase, or on loss landscape information like curvature or connectivity, a generator trained on our zoos could be trained to produce custom weights with predictable properties.
@@ Conclusion
Our work introduces a dataset of neural networks that systematically spans multiple phases of training across different architectures and domains. By providing controlled diversity in the weight space, we aim to accelerate research on weight space learning, a growing field that treats trained weights as a data modality. This dataset can help researchers analyze when and how different training phases arise, develop phase-aware training and downstream methods, and gain deeper insights into the fundamental properties of neural networks.
@@
KS, LM and DB are supported by the Swiss National Science Foundation research project grant 10001118 and SPRIND through the ModelFoundry project. YZ and YY are supported by DOE under Award Number DE-SC0025584 and funding from Dartmouth College.
@@ Appendix A Dataset documentation
Please find the dataset at https://phasetransitions.modelzoos.cc [^2] Back-up link: https://github.com/ModelZoos/PhaseTransitionModelZoo. The repository contains instructions on how to access the dataset, code to use it, code used to generate the model zoos and the loss landscape metrics, as well as other related information.
@@ A.2 Model Zoo Generation
We generated the dataset for common computer vision, language, and SciML tasks and architectures to show generalization and maximize applicability to the community. We fixed the load-temperature grids by exploring the boundary cases first and establishing the presence of phase transitions, then filling in more resolution. For the vision and language zoos, we chose batch size as a temperature-like parameter, as it’s a common hyperparameter to vary. For the load-like parameter, we vary the width of the models by changing the number of channels or the size of the internal token, respectively. For the SciML zoo, we adapt the scheme to the specifics of the domain. To vary the temperature, we chose to change the learning rate to keep the batch size and collocation points constant. To vary the load, we change the complexity of the domain by varying $\beta$ parameter of the physical system. We base the GPT-model implementation and training setup on NanoGPT Karpathy (2025) and use Chinchilla scaling laws to find the ideal number of training steps Hoffmann et al. (2022). The full list of model zoo hyperparameters is given in Tables 4,5 and 6. For vision zoos, we use Random Cropping, horizontal flipping, and random rotations. Training ViTs on CIFAR100 required stronger data augmentation to achieve competitive performance. Therefore, we have applied a combination of random cropping, random erasing, color jitter, and RandAugment (Cubuk et al., 2019). After the initial tuning of the grids, the training of the model zoos was done on 16 DGX H100 GPUS in 30 days. The computation of loss landscape metrics was performed on the same hardware in 14 days.
Table 3: Summary of the content of each model zoo included in our dataset.
Architecture Dataset # models Load-like param. Temperature-like param.
ResNet-18 SVHN 192 Model width Batch size
ResNet-18 CIFAR-10 192 Model width Batch size
ResNet-18 CIFAR-100 192 Model width Batch size
ResNet-18 TinyImagenet 192 Model width Batch size
ResNet-50 SVHN 192 Model width Batch size
ResNet-50 CIFAR-10 192 Model width Batch size
ResNet-50 CIFAR-100 192 Model width Batch size
ResNet-50 TinyImagenet 192 Model width Batch size
ViT CIFAR-10 147 Model width Batch size
ViT CIFAR-100 147 Model width Batch size
GPT-2 OpenWebText 264 model_dim Batch size
MLP (PINN) 1D Convection 700 beta Learning rate
@@
Learning Rate $0.0001$, $0.0005$, $0.001$, $0.005$, $0.01$, $0.05$, $0.1$, $0.5$, $1.0$, $1.5$
Convection coefficient $\beta$ 11, 22, 33, 44, 55, 66, 77, 88, 99, $10$, $15$, $30$, $50$, $70$
@@ A.3 Model zoo evaluation
In this section, we test the general validity of the trained models as representatives of real-world models in a structured dataset. An overview of the models at the end of training is given in Tables 7 and 8. The results confirm that models are trained to competitive performance for their respective sizes.More nuanced information on the distribution of model performance on the temperature-load grid is shown in Figures 6 through 16. Similar to previous work, the zoos show distinct low train-loss regions, with smaller embedded regions within that generalize well. Test performance generally improves with decreasing load (increasing width), with a distinct peak phase where temperature and load are low enough, but not too low. The generalization gap correspondingly shows a superposition of both patterns. Further applications or loss landscape metrics likewise show clear phase transitions. Remarkably, the phase layout and loss landscape metrics generalize across different domains, tasks, architectures, and datasets.
@@ A.5 Hosting, Licensing, and Maintenance Plan
The dataset is made publicly available and licensed under the Creative Commons Attribution 4.0 International license (CC-BY 4.0). It can be accessed at https://phasetransitions.modelzoos.cc, together with the corresponding code and instructions on how to use it.
The dataset is hosted by the University of St. Gallen and will be made available to the general public with detailed instructions and examples through the project’s public GitHub repository. The University of St. Gallen will ensure maintenance and long-term availability. We further plan on extending the dataset towards more architectures, tasks, and domains, and invite the community to engage.